OpenStudy (sonyalee77):

Students on a field trip at an amusement park were asked their grade in school. The table shows the results of the survey. Grade Number of Students 4th 45 5th 25 6th 60 7th 50 8th 60 9th 45 What percent of the students surveyed were in 6th grade? Round your answer to the nearest percent, if necessary. A. 60% B. 21% C. 6% D. 11%

OpenStudy (anonymous):

% of students in the 6th grade = (6th grade students/total number of students) * 100% = (60/45+25+60+50+60+45) x 100% = (60/285) x 100% = 21.0526... = 21 ( to the nearest digit)